The difference between the two positive numbers is equal to the double of the smaller number. if the product of the two numbers is 242, find the numbers.

Answer by fractalier(6550)

Call the numbers x and y. x is the larger.
From the facts of the problem, we have
x - y = 2y and
xy = 242
The first one becomes x = 3y and we can substitute this into the second one and get
3y*y = 242
y^2 = 242/3
y = 8.98
x = 3y = 26.94
I'm thinking there was a mistake in the numbers in the problem...